5^2x+1+5^2x-3+5^2x-4=155/625
Answer by stanbon(75887)

5^2x+1+5^2x-3+5^2x-4=155/625
--------------
Multiply thru by 5^4 to get:
5^(2x+5) + 5^(2x+1) + 5^(2x) = 155
----------
Factor out 5^(2x) to get:
5^(2x)[5^5 + 5^1 + 1] = 155
5^(2x)[3125+6] = 155
5^2x = 155/3131
Take the log of both sides
2x(log(5) = log(155)-log(3131)
2x = -1.30535/0.69897
x = -0.9339..
=====================
Cheers,
Stan H.